OpenStudy (anonymous):

It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Brian to build the car on his own.

OpenStudy (welshfella):

work out what fraction of the car they can build in 1 hour for John we let it be 1 / x ( letting the time he takes to build the car = x hours) for Brian it will be 1/ (x + 15) when they both work together , it takes 4 hours to build the car therefore they build 1/4 of the car in 1 hour. so we have the equation 1/x + 1/(x + 150 = 1/4 solve for x and you have x +15 will be the time that brian takes to build car on his own

OpenStudy (welshfella):

* thats 1/x + 1/(x+15) = 1/4

OpenStudy (welshfella):

can you solve that?

OpenStudy (anonymous):

x= -12,5 right?? @welshfella